# Spectral Graph Pruning Against Over-Squashing and Over-Smoothing

### v1

@@ Appendix E Additional Results
We perform semi-supervised node classification on the following datasets: Cora (McCallum et al., 2000), Citeseer (Sen et al., 2008) and Pubmed (Namata et al., 2012). We report results on Chameleon, Squirrel, Actor and the WebKB datasets consisting of Cornell, Wisconsin and Texas. Our baselines include GCN (Kipf & Welling, 2017) without any modifications to the original graph, DIGL by Gasteiger et al. (2019), SDRF by Topping et al. (2022), and FoSR by Karhadkar et al. (2023). We adopt the public implementations available and tune the hyperparameters to improve the performance if possible. Our results are presented in Table 10. We compare GCN with no edge modifications, GCN+DIGL, GCN+SDRF, GCN+FoSR, GCN+EldanDelete where we delete the edges, GCN+EldanAdd where we add the edges according to the criterion from Lemma 2.1 and ProxyAdd and ProxyDelete which use Equation (1) to optimize the spectral gap directly. The top performance is highlighted in bold. GCN+FoSR outperforms all methods on Cora, Citeseer and Pubmed, which are homophilic. Yet, GCN+ProxyAdd is more effective in increasing the spectral gap (see Table 15). On the remaining six datasets, our proposed methods both with edge deletions and additions outperform FoSR and SDRF, while outperforming all other baselines on all datasets. For training details and hyperparameters, please refer to the Appendix 18.
@@
GCN+FoSR 91.44±3.16 82.13±4.29 91.49±1.89 53.91±8.67 58.63±9.55 63.50±11.07 38.01±7.48 46.64±7.70 50.73±6.93
GCN+EldanDelete 87.60±0.18 78.68±0.54 87.33±0.07 65.13±13.02 67.84±7.65 70.53±6.70 43.65±9.88 52.51±8.12 48.89±7.89
GCN+EldanAdd 88.38±0.12 79.45 ±0.37 87.17±0.14 69.05±6.17 64.08±6.58 67.10±8.91 43.64±10.00 48.09±7.30 51.66±6.50
GCN+ProxyAdd 89.10±0.70 78.94±0.54 87.54±0.24 66.54±9.56 67.75±7.96 74.21±10.64 43.45±9.93 54.30±6.27 48.85±6.14
GCN+ProxyDelete 87.51±0.81 78.68 ±0.55 87.39±0.11 66.60 ± 6.50 66.36±7.17 72.36±7.88 43.52±9.64 55.88±5.48 48.90±7.68
@@ E.2 Graph Classification with GCN and R-GCN
We conduct experiments on graph classification with a GCN (Kipf & Welling, 2017) and R-GCN (Battaglia et al., 2018) backbone to demonstrate the effectiveness of our proposed rewiring algorithms. Our experimental setting is same as that of FoSR by Karhadkar et al. (2023), with the difference being we tune our hyperparameters on 10 random splits instead of 100. The final test accuracy is averaged over 5 random splits of data. We compare our results with FoSR by Karhadkar et al. (2023). For the IMDB-BINARY, REDDIT-BINARY and COLLAB datasets there are no node features available and have to be created. For fair comparison we run FoSR on these datasets. For ENZYMES and MUTAG the results are taken from the values reported in the paper. The results are reported in Table 11 and 12. From the tables it is clear that our proposed algorithms are effective in increasing the generalization performance even for graph classification tasks.
@@ Appendix F Update Period, Empirical Runtimes and Spectral Gap Comparisons
In §3.1, we have discussed the time complexity analysis of our proposed algorithms. Recall, our algorithm has a hyperparameter $M$, the number of edges to delete after ranking the edges using our proxy. For edge additions, the candidate edges that can be added are large, thus we can resort to sampling a constant set of edges to speed up the process. All of our experiments in §4 were conducted with $M=1$. However, it is possible to further reduce the overall runtimes by tuning the value of $M$, that is, how many edges can we modify before we have to recalculate the proxy to rank the edges again. This is shown in Table 14, where we compare our algorithms with $M=1$ and $M=10$, for 50 edge modifications. It is clear that although $M=1$ leads to better spectral gap improvement, $M=10$ is also a valid updating period which induces enough spectral gap change while simultaneously bringing down the runtime (also presented in Table 1) considerably, specially for large graphs. To further evaluate the trade-off between the update period and its effect on GNN test accuracy, we use ProxyAdd and ProxyDelete with different $M$ updates on Cora and Texas dataset to modify 50 and 20 edges respectively. This is shown in Figure 7. Although a more frequent update points to better test accuracy, update periods with $\{5,10\}$ also yield competitive results. Thus reinforcing the fact that, our proposed methods can be computationally efficient and can help in improving the generalization. In Table 15 we report the spectral gap changes induced by FoSR (Karhadkar et al., 2023), our proposed Eldan criterion based addition and deletions and also the Proxy versions of addition and deletions. In Table 16 we provide the runtimes for large heterophilic datasets (Platonov et al., 2023).
@@
Table 14: Empirical runtime (RT) comparisons with different update period for the criterion for 50 edges. We also report the spectral gap before (SG.B) and after rewiring (SG.A).
@@ Appendix G Pruning at Initialization for Graph Lottery Tickets
In Table 17, we present the results for Pruning at Initialization for finding graph lottery tickets. We first prune the input graph to required sparsity level and then the weights are iteratively pruned by magnitude similar to the approach proposed by (Chen et al., 2021). From the table it is clear that, at least for moderate graph sparsity (GS) levels for Cora dataset, that is around GS = $18.75\%$, our proposed EldanDelete-UGS and ProxyDelete attain comparable performance to UGS. On Pubmed for different graph sparsity levels we outperform UGS. While, our method fails to identify winning tickets for Citeseer. We use the public implementation by the authors (Chen et al., 2021) for all our lottery ticket experiments. For all experiments we report the test accuracy on node classification averaged over 3 runs. Except for Pubmed which could only be averaged over 2 runs.
@@
Appendix H Hyperparameters
We instantiate a 2-layered GCN (Kipf & Welling, 2017) for semi-supervised node classification, the Planetoid datasets (Cora, Citeseer and Pubmed) are available as pytorch geometric datasets. For the WebKB datasets we use the updated ones given by Platonov et al. (2023). We use a 60/20/20 split for training/testing/validation respectively for all datasets. We perform extensive hyperparameter tuning on the validation set and finally report test accuracy averaged over 10 random splits of the data. We use the largest connected component wherever available. The same experimental settings hold for other baselines DIGL, SDRF and FoSR. For node classification using R-GCNs, we also use a 3 layered GCN, this is highlighted in Table 22 with other hyperparameters. For graph classification, we use the same experimental setup as (Karhadkar et al., 2023), we use a 4-layered GCN and R-GCN versions. For the larger heterophilic datasets, we use the experimental setup given by the authors (Platonov et al., 2023). We set the learning rate to $43e-$ 4, dropout to $0.32$, the hidden dimension size to $512$. For GATs, the attention heads are set to $8$. The datasets are split into 50%/25%/25% for train, test and validation respectively. We tune our edge modification algorithms on the validation set. The final test accuracy is reported averaged over 10 random splits run for 1000 steps. We use PyTorch Geometric and DGL library for our experiments.
@@ Appendix H Hyperparameters
Table 22: Hyperparameters for R-GCN+ProxyAdd on Node Classification
Dataset LR Layers HiddenDimension Dropout ProxyAdd
Cora 0.01 2 32 0.3130296 50
Citeseer 0.01 2 64 0.3130296 250
Pubmed 0.01 2 32 0.4130296 100
Cornell 0.001 3 128 0.3130296 05
Wisconsin 0.001 3 128 0.3130296 25
Texas 0.01 3 128 0.3130296 20
Actor 0.001 3 128 0.5130296 25
Chameleon 0.001 3 128 0.4130296 100
Squirrel 0.001 3 128 0.3130296 5
Dataset LR Dropout
Hidden
Dimension
EldanAdd
ENZYMES 0.001 0.2130296 32 20
MUTAG 0.001 0.3130296 32 20
IMDB-BINARY 0.001 0.3130296 32 10
REDDIT-BINARY 0.001 0.2130296 32 10
COLLAB 0.001 0.21 32 10
PROTEINS 0.001 0.3130296 32 10
Dataset LR Dropout
Hidden
Dimension
ProxyAdd
ENZYMES 0.001 0.2130296 32 20
MUTAG 0.001 0.3130296 32 20
IMDB-BINARY 0.001 0.3130296 32 10
REDDIT-BINARY 0.001 0.2130296 32 10
COLLAB 0.001 0.21 32 10
PROTEINS 0.001 0.3130296 32 10

### v2

@@ Appendix E Additional results
We perform semi-supervised node classification on the following datasets: Cora (McCallum et al., 2000), Citeseer (Sen et al., 2008) and Pubmed (Namata et al., 2012). We report results on Chameleon, Squirrel, Actor and the WebKB datasets consisting of Cornell, Wisconsin and Texas. Our baselines include GCN (Kipf & Welling, 2017) without any modifications to the original graph, DIGL by Gasteiger et al. (2019), SDRF by Topping et al. (2022), and FoSR by Karhadkar et al. (2023). We adopt the public implementations available and tune the hyperparameters to improve the performance if possible. Our results are presented in Table 10. We compare GCN with no edge modifications, GCN+DIGL, GCN+SDRF, GCN+FoSR, GCN+RandomDelete, GCN+EldanDelete where we delete the edges, GCN+EldanAdd where we add the edges according to the criterion from Lemma 3.1 and ProxyAdd and ProxyDelete which use Equation (1) to optimize the spectral gap directly. The results for GCN+BORF (Nguyen et al., 2023) are taken from the paper directly, hence NA for some datasets. The top performance is highlighted in bold. GCN+FoSR outperforms all methods on Cora, Citeseer and Pubmed, which are homophilic. Yet, GCN+ProxyAdd is more effective in increasing the spectral gap (see Table 16). On the remaining six datasets, our proposed methods both with edge deletions and additions outperform FoSR and SDRF, while outperforming all other baselines on all datasets. For training details and hyperparameters, please refer to the Appendix 19.
@@
GCN+FoSR 91.44±0.39 82.13±0.31 91.49±0.10 53.91±1.47 58.63±1.46 63.50±1.75 38.01±0.21 46.64±0.63 50.73±0.37
GCN+EldanDelete 87.60±0.18 78.68±0.54 87.33±0.07 65.13±1.50 67.84±1.45 70.53±1.23 43.65±0.21 52.51±0.55 48.89±0.40
GCN+EldanAdd 88.38±0.12 79.45 ±0.37 87.17±0.14 69.05±1.50 64.08±1.63 67.10±1.13 43.64±0.25 48.09±0.59 51.66±0.45
GCN+ProxyAdd 89.10±0.70 78.94±0.54 87.54±0.24 66.54±1.41 67.75±1.64 74.21±1.25 43.45±0.20 54.30±0.59 48.85±0.39
GCN+ProxyDelete 87.51±0.81 78.68 ±0.55 87.39±0.11 66.60 ± 1.67 66.36±1.33 72.36±1.35 43.52±0.22 55.88±0.70 48.90±0.39
GCN+RandomDelete 87.30±0.31 78.34±0.38 87.15±0.16 63.97±2.50 61.71±2.73 63.97±5.41 29.57±0.44 44.07±1.04 40.63±0.41
GCN+BORF 87.50±0.20 73.80±0.20 NA 50.80±1.11 50.30±0.90 49.40±1.20 NA 61.50±0.40 NA
@@ E.2 Graph classification with GCN and R-GCN
We conduct experiments on graph classification with a GCN (Kipf & Welling, 2017) and R-GCN (Battaglia et al., 2018) backbone to demonstrate the effectiveness of our proposed rewiring algorithms. Our experimental setting is the same as that of FoSR by Karhadkar et al. (2023), with the difference being we tune our hyperparameters on 10 random splits instead of 100. The final test accuracy is averaged over 5 random splits of data. We compare our results with FoSR by Karhadkar et al. (2023). For the IMDB-BINARY, REDDIT-BINARY and COLLAB datasets there are no node features available and have to be created. For fair comparison we run FoSR on these datasets. For ENZYMES and MUTAG the results are taken from the values reported in the paper. The results are reported in Table 11 and 12. From the tables it is clear that our proposed algorithms are effective in increasing the generalization performance even for graph classification tasks.
@@ Appendix F Update period, empirical runtimes and spectral gap comparisons
In §4.1, we have discussed the time complexity analysis of our proposed algorithms. Recall, that our algorithm has a hyperparameter $M$, the number of edges to delete after ranking the edges using our proxy. For edge additions, the candidate edges that can be added are large, thus we can resort to sampling a constant set of edges to speed up the process. All of our experiments in §1 were conducted with $M=1$. However, it is possible to further reduce the overall runtimes by tuning the value of $M$, that is, how many edges we can modify before we have to recalculate the proxy to rank the edges again. This is shown in Table 15, where we compare our algorithms with $M=1$ and $M=10$, for 50 edge modifications. It is clear that although $M=1$ leads to better spectral gap improvement, $M=10$ is also a valid updating period which induces enough spectral gap change while simultaneously bringing down the runtime (also presented in Table 14) considerably, especially for large graphs. To further evaluate the trade-off between the update period and its effect on GNN test accuracy, we use ProxyAdd and ProxyDelete with different $M$ updates on Cora and Texas datasets to modify 50 and 20 edges respectively. This is shown in Figure 8. Although a more frequent update points to better test accuracy, update periods with $\{5,10\}$ also yield competitive results. Thus reinforcing the fact that our proposed methods can be computationally efficient and can help in improving the generalization. In Table 16 we report the spectral gap changes induced by FoSR (Karhadkar et al., 2023), our proposed Eldan criterion based addition and deletions and also the Proxy versions of addition and deletions. In Table 17 we provide the runtimes for large heterophilic datasets (Platonov et al., 2023).
@@
Table 14: Runtimes for 50 edge modifications in seconds.
Method Cora Citeseer Chameleon Squirrel
FoSR 4.69 5.33 5.04 19.48
SDRF 19.63 173.92 17.93 155.95
ProxyAdd 4.30 3.13 1.15 9.12
ProxyDelete 1.18 0.86 1.46 7.26
Table 15: Empirical runtime (RT) comparisons with different update periods for the criterion for 50 edges. We also report the spectral gap before (SG.B) and after rewiring (SG.A).
@@ Appendix G Pruning at initialization for graph lottery tickets
In Table 18, we present the results for Pruning at Initialization for finding graph lottery tickets. We first prune the input graph to the required sparsity level and then the weights are iteratively pruned by magnitude similar to the approach proposed by (Chen et al., 2021). From the table it is clear that, at least for moderate graph sparsity (GS) levels for Cora dataset, that is around GS = $18.75\%$, our proposed EldanDelete-UGS and ProxyDelete attain comparable performance to UGS. On Pubmed for different graph sparsity levels we outperform UGS. Meanwhile, our method fails to identify winning tickets for Citeseer. We use the public implementation by the authors (Chen et al., 2021) for all our lottery ticket experiments. For all experiments we report the test accuracy on node classification averaged over 3 runs. Except for Pubmed which could only be averaged over 2 runs.
@@
Appendix H Training details and hyperparameters
We instantiate a 2-layered GCN (Kipf & Welling, 2017) for semi-supervised node classification, the Planetoid datasets (Cora, Citeseer and Pubmed) are available as pytorch geometric datasets. For the WebKB datasets we use the updated ones given by Platonov et al. (2023). We use a 60/20/20 split for training/testing/validation respectively for all datasets. We perform extensive hyperparameter tuning on the validation set and finally report test accuracy averaged over 10 splits of the data (Chen et al., 2018). We use the largest connected component wherever available. The same experimental settings hold for other baselines DIGL, SDRF and FoSR. For node classification using R-GCNs, we also use a 3 layered GCN, this is highlighted in Table 20 with other hyperparameters. For graph classification, we use the same experimental setup as (Karhadkar et al., 2023), we use a 4-layered GCN and R-GCN versions. For the larger heterophilic datasets, we use the experimental setup given by the authors (Platonov et al., 2023). We set the learning rate to $\{3e-3,3e-4\}$, dropout to $0.32$, and the hidden dimension size to $512$. For GATs, the attention heads are set to $8$. The datasets are split into 50%/25%/25% for train, test and validation respectively. We tune our edge modification algorithms on the validation set. The final test accuracy is reported as averaged over 10 random splits run for 1000 steps. Skip connections and normalization (Ba et al., 2016) is used to facilitate training deeper models. We use PyTorch Geometric and DGL library for our experiments. All experiments were done on 2 V100 GPUs. Our code https://github.com/RelationalML/SpectralPruningBraess is available.
@@ Appendix H Training details and hyperparameters
Table 20: Hyperparameters for R-GCN+ProxyAdd on node classification
Dataset LR Layers HiddenDimension Dropout ProxyAdd
Cora 0.01 2 32 0.3130296 50
Citeseer 0.01 2 64 0.3130296 250
Pubmed 0.01 2 32 0.4130296 100
Cornell 0.001 3 128 0.3130296 05
Wisconsin 0.001 3 128 0.3130296 25
Texas 0.01 3 128 0.3130296 20
Actor 0.001 3 128 0.5130296 25
Chameleon 0.001 3 128 0.4130296 100
Squirrel 0.001 3 128 0.3130296 5
Dataset LR Dropout
Hidden
Dimension
EldanAdd
ENZYMES 0.001 0.2130296 32 20
MUTAG 0.001 0.3130296 32 20
IMDB-BINARY 0.001 0.3130296 32 10
REDDIT-BINARY 0.001 0.2130296 32 10
COLLAB 0.001 0.21 32 10
PROTEINS 0.001 0.3130296 32 10
Dataset LR Dropout
Hidden
Dimension
ProxyAdd
ENZYMES 0.001 0.2130296 32 20
MUTAG 0.001 0.3130296 32 20
IMDB-BINARY 0.001 0.3130296 32 10
REDDIT-BINARY 0.001 0.2130296 32 10
COLLAB 0.001 0.21 32 10
PROTEINS 0.001 0.3130296 32 10